OpenStudy (anonymous):

How long does it take to drive from Minneapolis to Chicago through Green Bay? Route Hours to Travel by Car Minneapolis to 4 5/6 Green Bay Green Bay to Chicago 3 2/3

OpenStudy (anonymous):

1 1/6 7 7/16 8 1/2 8 2/3

OpenStudy (anonymous):

PLEASE HELP!!

OpenStudy (anonymous):

it would be 8 1/2

OpenStudy (anonymous):

4 5/6 + 3 4/6 =7 9/6 = 8 3/6 = 8 1/2

OpenStudy (anonymous):

ok thank you!!